How much of a 12% saline solution should be mixed with a 4% saline solution to obtain 40 milliliters of a 6% saline solution? Use a system of equations and the elimination method to solve the problem.
Answer by stanbon(75887)

How much of a 12% saline solution should be mixed with a 4% saline solution to obtain 40 milliliters of a 6% saline solution? Use a system of equations and the elimination method to solve the problem.
-----
Quantity:: t + f = 40 ml
Active:::0.12t+0.04f = 0.06*40
------------------------------------
Modify for elimination:
4t + 4f = 4*40
12t + 4f = 6*40
------------------------
Subtract and solve for "t"::
8t = 2*40
t = 10 ml (amt of 12% solution needed)
------------
Cheers,
Stan H.
------------